2. Suppose you have a four-sided dice, a six-sided dice and a 12-sided dice, each dice is fair. You roll the three dice once and let Z be the number of 4s showing. Find E(Z) using the indicator method.
Answer #1

let X1 ; X2 and X3 are binomial indicator variable which can take values 1 if shows a four and 0 if does not show a 4.

hence Z =X1+X2+X3

therefore E(Z)=E(X1)+E(X2)+E(X3)=p1+p2+p3 =(1/4)+(1/6)+(1/12)=0.5